You make a one-time deposit of $3600 into an account earning 6% interest compounded quarterly. How much will you have in the account in 8 years? Round to the nearest cent.
Answer #1

Amount deposited = PV = $3600

Annual interest rate = r = 6%

It is compounded quarterly

Qaurterly interest rate = rq = 6%/4 = 1.5%

No. of quarters in 8 years = n = 8*4 = 32

Future value is calculated using the formula:

Future value after 8 years = PV*(1+rq)n = 3600*(1+1.5%)32 = 5797.16755255867 ~ $5797.17 (Rounded to nearest cent)

Answer -> 5797.17